A man sold 10 eggs for Rs. 36 and thus gained 20% profit. At what price did he buy an
egg?

Answer by checkley79(3341)

36/10=3.6 selling price per egg.
1.2x=3.6
x=3.6/1.2
x=3 cost of each egg bought.
Proof:
3*1.2=3.6
3.6=3.6